OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, –3) and has a slope of one-half?

OpenStudy (anonymous):

\[\left( y - y1 \right) = m \times \left( x - x1 \right)\]

hero (hero):

y = mx + b -3 = .5(-2) + b Solve for b

OpenStudy (anonymous):

y - y1 = m(x - x1) y - - 3 = 1/2(x - - 2) y + 3 = 1/2x + 1 y = 1/2x + 1 - 3 y = 1/2x - 2. Now change it into standard form 2y = x - 4 x - 2y = 4